OpenStudy (praxer):

If 1000! is divisible by 13^n than find the largest value of n.

OpenStudy (praxer):

@ganeshie8

OpenStudy (praxer):

@ikram002p bachao bachao.. :)

OpenStudy (praxer):

I wasted last one hour in posting the questions. OS Seems damn sloppy today..

OpenStudy (praxer):

my queries never bump.. :(

OpenStudy (ikram002p):

well (100!) the idea is to get all numbers that devieds 13 which are less than 1000 like :- 13,2.13,3.13..... n.13 (PS dot means x ) 13^2,2.13^2 , .., n.13^2 13^3,2.13^3,.....n.13^3 ...... 13^n,2.13^n,.....n13^n and count them . do you know legendre theorem ? (its already been given in this formula ) http://www.cut-the-knot.org/blue/LegendresTheorem.shtml

OpenStudy (ikram002p):

OS is lagging :)

OpenStudy (ikram002p):

so according to the link i gave greatest n = 1000-(sum of digits of (1000)base 13) /13-1

OpenStudy (zarkon):

\[\left\lfloor{\frac{1000}{13}}\right\rfloor+\left\lfloor{\frac{1000}{13^2}}\right\rfloor\]

OpenStudy (ikram002p):

Zacron , but following this is much easier greatest n = 1000-(sum of digits of (1000)base 13) /13-1 according to this rule http://prntscr.com/4s4x97

OpenStudy (ikram002p):

thus greatest n = (1000-(sum digits of (5bc)))/(13-1) =1000-28/12=81 thus 13^81 | 1000!

OpenStudy (ikram002p):

yeahhh so i was correct ,first time i used legendre lol